# Post-detection inference for sequential changepoint localization

### v2

@@ 15            $\mathcal{C}=\mathcal{C}\cup\{t\}$
For any fixed $\alpha\in(0,1)$, $B,N\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$, $T\in\mathbb{N}$ and change detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}$ defined in (2.13) satisfies
$\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-{\alpha}.$ (2.14)
Proof.
First note the fact that $\mathbb{P}_{T}(\tau<T)=\mathbb{\mathbb{P}}_{\infty}(\tau<T),$ since the event $[\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
Step 1 [Relate the level of testing ${H}_{0,T}$ and the conditional coverage]: Defining $C_{t,L}=\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})$, the level of testing ${H}_{0,T}$ is $P_{T}(M_{T}>C_{t,L}).$ Thus,
$\mathbb{P}_{T}(M_{T}\leq C_{t,L})$
$=\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\})$
$=\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\},\tau\geq T)+\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\},\tau<T)$
$\leq\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq C_{t,L}\}\mid\tau\geq T)\mathbb{P}_{T}(\tau\geq T)+\mathbb{\mathbb{P}}_{T}(\tau<T)$
$=\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)+1-\mathbb{\mathbb{P}}_{\infty}(\tau\geq T),$
where we twice invoked $\mathbb{P}_{T}(\tau<T)=\mathbb{\mathbb{P}}_{\infty}(\tau<T)$ as explained earlier. So, we obtain
$\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{P_{T}(M_{T}>C_{t,L})}{P_{\infty}(\tau\geq T)}.$ (2.15)
Step 2 [Establish that our test for each ${H}_{0,t}$ has level $\alpha\times\mathbb{P}_{\infty}(\tau\geq t)$]: First, observe that $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$ are i.i.d. under $H_{0,t}$. Hence, $M_{t}\leq\operatorname{Quantile}(1-c;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B})$ $\iff M_{t}$ is one among the $\lceil(1-c)(B+1)\rceil$ smallest of $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$. So, for any fixed $c\in(0,1)$,
$\mathbb{P}_{t}(M_{t}\leq\operatorname{Quantile}(1-c;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}))\geq\frac{\lceil(1-c)(B+1)\rceil}{B+1}\geq 1-c\text{ for all }t.$
Since $\hat{\alpha}_{t}$ is independent of $M_{t}$, it follows that
$\mathbb{P}_{t}(M_{t}\leq C_{t,\infty}\mid\hat{\alpha}_{t})\geq\frac{\lceil(1-\hat{\alpha}_{t})(B+1)\rceil}{B+1}\geq 1-\hat{\alpha}_{t},\text{ almost surelyfor all }t.$
Therefore, we have
$\mathbb{P}_{t}(M_{t}\leq C_{t,\infty})=\mathbb{E}_{t}(\mathbb{P}_{t}(M_{t}\leq C_{t,\infty}\mid\hat{\alpha}_{t}))\geq\mathbb{E}_{t}(1-\hat{\alpha}_{t})=1-\alpha\mathbb{P}_{t}(\tau\geq t).$
As noted earlier, $\mathbb{P}_{t}(\tau\geq t)=\mathbb{P}_{\infty}(\tau\geq t)$ and hence, $\mathbb{P}_{t}(M_{t}\leq C_{t,\infty})\geq 1-\alpha\mathbb{P}_{\infty}(\tau\geq t).$ Also, for all $L\in\mathbb{N}$, it follows from the definitions that $M_{t,L}^{j}\geq M_{t,\infty}^{j}$. Therefore, $C_{t,L}\geq C_{t,\infty}$. Thus, for all $t$,
$\mathbb{P}_{t}(M_{t}>C_{t,L})\leq\mathbb{P}_{t}(M_{t}>C_{t,\infty})\leq\alpha\mathbb{P}_{\infty}(\tau\geq t)$ (2.16)
Therefore, combining (2.15) and (2.16), we obtain $\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha.$

### v3

@@ 5 Confidence set for pre- and post-change parameters
Let $\mathcal{C}$ be a confidence set, e.g., (3.8) or (4.5), such that for any $\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1},T\in\mathbb{N}$, change detection algorithm $\mathcal{A}$ satisfying Assumption 3.3 and $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$, we have: $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha$. Then, for any $T\in\mathbb{N}$,
$\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{0}\in\displaystyle\cup_{t\in\mathcal{C}}\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*})\mid\tau\geq T\right)\geq 1-\alpha-\eta_{0},\text{ and }$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\in\displaystyle\cup_{t\in\mathcal{C}}\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*})\mid\tau\geq T\right)\geq 1-\alpha-\eta_{1}.$
6 Experiments
We validate our methods across various settings and demonstrate their practicality.
Known pre- and post-change (Setting I): Here, we draw $X_{1},\cdots,X_{T-1}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ and $X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}N(1,1)$, with $T$ being the only unknown parameter. We employ the CUSUM detector, $\tau=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\prod_{i=j}^{n}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ with $A=1000$. After detection, we construct both the proposed confidence sets (universal (3.6) and adaptive (4.1)), with $B=N=100$, $\alpha=0.05$. Fig. 6.1 visualizes the confidence sets (4.1) and point estimates (3.1) across 555 runs. Table 6.2 has the (conditional) average size and coverages of the confidence sets, the (conditional) average absolute deviation of $\hat{T}$, and delay of detection, i.e, $\tau-T$, across 500500500 runs.
Figure 6.1: Setting I: The first $T-1$ observations are drawn from $N(0,1)$ and the rest from $N(1,1)$. The point estimates (3.1) are shown in a vertical red dashed line, and confidence sets (adaptive (4.1)) are shown in red points, with $B=N=100$, $\alpha=0.05,L=\infty$. Results of 555 independent simulations are shown.
Table 6.2: Setting I: Known pre- and post-change.
$T$ A Conditional Coverage Conditional Size
$|\hat{T}-T|$
Delay
Universal Adaptive Universal Adaptive (Conditional) (Conditional)
100 1000 0.98 0.95 15.63 12.34 2.85 13.97
500 1000 0.98 0.95 15.77 12.57 2.62 13.22
Known pre- and parametric post-change (Setting II): We perform experiments with $F_{0}=N(0,1)$ (known pre-change), post-change class $\mathcal{P}_{1}=\{N(\theta,1):\theta\in\Theta_{1}\}$, ; we vary $\Theta_{1}$ and $T$. The post-change samples are drawn from $N(1,1)$ (unknown). For detection, we use a weighted CUSUM detector:
$\tau_{\text{wcusum}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (6.1)
with $A=1000$, a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$, $d=0.2$, $\theta=0.75$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>\theta\}$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized mixture over $\mathcal{P}_{1}$. We use
$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\theta_{0}}(X_{i})},$ (6.2)
where $\hat{\theta}_{1,j:n}$ is the MLE of $\theta_{1}\in\Theta_{1}$ based on $X_{j},\cdots,X_{n}$. After detection, we construct both the proposed confidence sets for $T$ (universal (3.6) with $\alpha=0.075$ and adaptive (4.4) [^6] Implementation details of the adaptive method in Setting II (and Setting III) are the same as discussed in Section B.1 (and Section C.1) of the Supplementary. with $\alpha=0.05,\beta=0.025$) with target conditional coverage 0.9250.9250.925. For implementing (3.6) , we use $R_{n}^{(t)}=\prod_{i=t}^{n}\frac{f_{\theta_{0}}(X_{i})}{f_{\theta_{1}^{*}}(X_{i})},$ where ${\theta_{1}^{*}}$ is the closest element of $\Theta_{0}$ from $\Theta_{1}$, and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}w(\theta)d\theta$, with $w$ being the same weight mentioned above. In Table 6.4, we report the (conditional) average size and coverages, absolute deviation of $\hat{T}$ , and detection delay $\tau-T$ across 500500500 runs. Fig. 6.3 visualizes the confidence sets (3.6) and point estimates (B.2) across 555 random runs.
Figure 6.3: Setting II: First $T-1$ samples are drawn from $N(0,1)$ and the rest from $N(1,1)$. The point estimates (B.2) are shown in a vertical red dashed line, and confidence sets (universal (3.6)) are shown in red points. $F_{0}=N(0,1)$, $\mathcal{P}_{1}=\{N(\mu,1):\mu\geq 0.9\}$, $N=100$, $\alpha=0.075$. Results of 555 independent simulations are shown.
Table 6.4: Known pre-change but unknown post-change. The pre-change parameter $\theta_{0}=0$ is known, the true (unknown) post-change parameter is $\theta_{1}=1$, while the algorithm only knows $\theta_{1}\in\Theta_{1}$, which we mention in the second column. (Setting II)
$T$ $\Theta_{1}$ Conditional Coverage Conditional Size
$|\hat{T}-T|$
Delay
Universal Adaptive Universal Adaptive (Conditional) (Conditional)
100 [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.98 0.97 22.21 25.72 3.95 16.87
100 [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.97 0.97 17.85 24.11 3.67 16.21
500 [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.98 0.97 22.89 30.83 3.45 17.63
500 [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.97 0.98 18.14 28.56 3.48 15.81
Composite parametric pre- and post-change (Setting III): Here, we draw $X_{1},\cdots,X_{T-1}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ and $X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}N(1,1)$. Now we conduct experiments for pre-change class $\mathcal{P}_{0}=\{N(\theta,1):\theta\in\Theta_{0}\}$, post-change class $\mathcal{P}_{1}=\{N(\theta,1):\theta\in\Theta_{1}\}$; we vary $\Theta_{0},\Theta_{1}$ and $T$. For detection, we employ a weighted CUSUM-type detector
$\tau_{\text{wcs-ripr}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta^{*}_{0}}(X_{i})}dW(\theta)\geq A\right\},$ (6.3)
with $A=1000$, ${\theta^{*}_{0}}$ is the “closest” element of $\Theta_{0}$ from $\Theta_{1}$, and for the post-change parameter, the same discrete weight distribution as in setting II is considered. We define $\hat{\theta}_{0,1:j}$ to be the MLE of $\theta_{0}\in\Theta_{0}$ based on data $X_{1},\cdots,X_{j}$. We use
$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\hat{\theta}_{0,1:j-1}}(X_{i})},$ (6.4)
where $\hat{\theta}_{1,j:n}$ is the MLE of $\theta_{1}\in\Theta_{1}$ based on $X_{j},\cdots,X_{n}$ and $\hat{\theta}_{0,:j-1}$ is the MLE of $\theta_{0}\in\Theta_{0}$ based on $X_{1},\cdots,X_{j-1}$. After detection, we construct both the proposed confidence sets (universal (3.8) with $\alpha=0.1$ and adaptive (4.5) with $\alpha=0.05,\beta=\gamma=0.025$) for $T$, with $N=B=100$, both having target conditional coverage 0.90.90.9. For implementing (3.8), we use $R_{n}^{(t)}=\int_{\theta\in\Theta_{0}}\prod_{i=t}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{1}^{*}}(X_{i})}w^{\prime}(\theta)d\theta,$ with $w^{\prime}(\theta)$ being the discrete distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta-d,\theta_{3}=\theta-2d,\cdots,\theta_{10}=\theta-9d$, $d=0.2$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>\theta\}$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$ and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}^{*}}(X_{i})}w(\theta)d\theta$, with $w(\theta)$ being the same as mentioned in Setting II. $\theta_{i}^{*}$ is the closest element of $\Theta_{i}$ from the other space. Fig. 6.5 visualizes the confidence sets (3.8) and point estimates (C.2) across 5 random runs. Table 6.6 reports the results averaged across 500500500 runs.
Figure 6.5: Setting III: First $T-1$ samples are drawn from $N(0,1)$ and the remaining samples from $N(1,1)$. The point estimates (C.2) are shown in a vertical red dashed line. Confidence sets (universal (3.8)) are shown in red points. $\mathcal{P}_{0}=\{N(\mu,1):\mu\leq 0.1\}$, $\mathcal{P}_{1}=\{N(\mu,1):\mu>0.9\}$, $N=100$, $\alpha=0.1$. Results of 555 random simulations are shown.
Table 6.6: Confidence interval for $T$ (Setting III): unknown pre- and post-change. The true (unknown) values of the parameters are $\theta_{0}=0$, $\theta_{1}=1$, while the algorithm only knows $\theta_{i}\in\Theta_{i},i=0,1$, which are mentioned in the second and third columns.
$T$ $\Theta_{0}$ $\Theta_{1}$ Conditional Coverage Conditional Size
$|\hat{T}-T|$
Delay
Universal Adaptive Universal Adaptive (Cond.) (Cond.)
100 (,0.25]($\infty$,0.25]( ∞ , 0.25 ] [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.98 0.97 26.91 26.84 4.36 25.81
100 (,0.1]($\infty$,0.1]( ∞ , 0.1 ] [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.97 0.97 18.63 25.71 4.19 23.15
500 (,0.25]($\infty$,0.25]( ∞ , 0.25 ] [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.98 0.97 26.12 31.45 4.03 24.48
500 (,0.1]($\infty$,0.1]( ∞ , 0.1 ] [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.97 0.98 18.86 30.66 4.07 23.07
Known pre- and nonparametric post-change (Setting IV): Here, the pre-change samples are from known $F_{0}=\text{U}[0,1]$ (uniform distribution on [0,1][0,1][ 0 , 1 ]) and let the post-change class be $\mathcal{P}_{1}:=\left\{F\in\mathcal{M}:\mathbb{E}_{F}[X]\leq\mu\right\},$ where $\mathcal{M}$ is the set of all probability measures on $([0,1],\mathcal{B}([0,1]))$, with $\mu=0.25$. For detection, we use an e-detector [26] $\tau_{e}=\inf\left\{n\in\mathbb{N}:\max_{1\leq t\leq n}\prod_{i=t}^{n}\frac{\hat{p}_{i-1}^{\prime}(X_{i})}{{p}_{0}(X_{i})}\geq A\right\},$ where $\hat{p}_{i-1}^{\prime}(x)=\frac{1+\sum_{k=1}^{10}\mathds{1}(x\in B_{k})\sum_{j=t}^{i-1}\mathds{1}(X_{j}\in B_{k})}{9+t-i},$ for $x\in[0,1]$, with $A=1000$. Post-detection, we construct the confidence sets (3.6). The forward $t$-delay e-process $\{R^{(t)}_{n}\}_{n\geq t}$, we use numeraire e-variable for $\mathcal{P}_{1}$ against $F_{0}$, as in Section 5.2 of [15] to get $R^{(t)}_{n}=\prod_{i=t}^{n}(1+\lambda^{*}(X_{i}-\mu)),$ where $\lambda^{*}\in(0,1/\mu)$ is the solution of the equation $e^{\lambda^{*}}=\frac{1+\lambda^{*}(1-\mu)}{1-\lambda^{*}\mu}.$ For the backward e-processes $\{S^{(t)}_{n}\}_{n<t}$, we adopt the predictable plug-in method [19]. We divide the interval [0,1][0,1][ 0 , 1 ] into 101010 equal-sized bins $B_{i}=[\frac{i-1}{10},\frac{i}{10}),i=1,\cdots,10$ and perform histogram-based estimation. We construct $S^{(t)}_{n}=\prod_{i=n}^{t-1}\frac{\hat{p}_{i-1}(X_{i})}{{p}_{0}(X_{i})},$ where ${p}_{0}(x)=\frac{1}{10},$ and $\hat{p}_{i-1}(x)=\frac{1+\sum_{k=1}^{10}\mathds{1}(x\in B_{k})\sum_{j=i-1}^{t-1}\mathds{1}(X_{j}\in B_{k})}{9+t-i},x\in[0,1]$. We use the point estimator
$\hat{T}=\operatorname*{arg\,max}_{1\leq t\leq\tau_{e}}\prod_{i=t}^{\tau_{e}}\frac{\hat{p}_{i-1}^{\prime}(X_{i})}{{p}_{0}(X_{i})}.$ (6.5)
Table 6.7 contains average results across 500500500 runs and Fig. 6.8 visualizes across 555 runs.
[…]
T
True (unknown) post-
change density, $f_{1}(x)$
Coverage
Coverage
Size
$|\hat{T}-T|$
Delay
100 $4(1-x)^{3}$ 0.98 0.97 24.14 4.06 20.65
[…]
500 $4(1-x)^{3}$ 0.98 0.94 24.63 4.11 19.57
[…]
Figure 6.8: Setting IV: First $T-1$ samples are drawn from $U(0,1)$ and the remaining observations from the density $f(x)=4(1-x)^{3},x\in[0,1]$. The point estimate (6.5) is shown in a vertical red dashed line, and the confidence set (3.8) is shown in red points. Results of 555 random simulations are shown.
[…]
$\tau_{e}^{\prime}=\inf\left\{n\in\mathbb{N}:\max_{1\leq t\leq n}\prod_{i=t}^{n}e^{\nu_{i}X_{i}-\nu_{i}^{2}/2}\geq A\right\},$
with $A=1000$, $\nu_{i}=\min\{0,\frac{1}{i-1}\sum_{j=1}^{i-1}X_{j}\}-0.5.$ This $\mathcal{P}_{0}$ and $\tau_{e}^{\prime}$ satisfies Assumption 3.3 with $F_{0}^{*}=\text{N}(0.5,1)$. After detection, we construct the confidence sets (3.8). The forward $t$-delay e-process we use is $\{R^{(t)}_{n}\}_{n\geq t}=\prod_{i=t}^{n}e^{\lambda_{i}X_{i}-\lambda_{i}^{2}/2}$, where $\lambda_{i}=\max\{0.5,\frac{1}{i-t}\sum_{j=t}^{i-1}X_{i}\}$ and the backward e-processes is constructed similarly: $\{S^{(t)}_{n}\}_{n<t}=\prod_{i=n}^{t}e^{\mu_{i}X_{i}-\mu_{i}^{2}/2}$, where $\mu_{i}=\min\{0,\frac{1}{t-i}\sum_{j=i+1}^{t}X_{j}\}-0.5$ (see, e.g., Ch. 6.7.5 of [21]). We use the following point estimator:
$\hat{T}=\operatorname*{arg\,max}_{1\leq t\leq\tau_{e}}\prod_{i=t}^{\tau_{e}^{\prime}}e^{\nu_{i}X_{i}-\nu_{i}^{2}/2}.$ (6.6)
Figure 6.9: Setting V. First $T-1$ samples are drawn from $N(1,1)$ and the remaining samples are from $U(-1.2,0.8)$. Vertical red dashed lines show the point estimates (6.6). The confidence set (3.8) is marked in red points. Results of 555 independent simulations are shown.
Table 6.10: Setting V: Parametric pre-change, nonparametric post-change.
T
Coverage
Coverage
Size
$|\hat{T}-T|$
Delay
100 0.98 0.97 27.65 3.08 27.34
500 0.98 0.94 28.03 3.15 25.12
Comparison with [36]: We now compare our method with Ch. 3.5 of [36] on the Gaussian mean-change, having the (unknown) changepoint at $T=100$ or 500500500. We recall that, unlike our general nonasymptotic guarantees for any $\mathcal{A}$, Wu’s method only has asymptotic approximate guarantees in specific parametric settings, and only for CUSUM procedure [18], $\tau^{\prime}_{\text{cusum}}:=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}T_{j}\geq d\right\},$ where $T_{n}=\max\{0,T_{n-1}+X_{n}\}$ and $d$ is the threshold. We used $d=8.59$ and 7.567.567.56 for ${N}(-0.25,1)$ vs. ${N}(0.25,1)$ and ${N}(-0.3,1)$ vs. ${N}(0.3,1)$ respectively. We implement Wu’s confidence set of the form $V_{c}\cup[L_{s},\hat{\nu})$, where $\hat{\nu}=\max\{k<\tau:T_{k}=0\}$, $V_{c}=\{k\geq\hat{\nu}:T_{k}\leq c\}$, and $L_{k}$ denote the $k$-th last zero point of $T_{n}$ counted backward starting from $L_{0}:=\hat{\nu}$, with their proposed approximations (for the case when $\theta_{1}=-\theta_{0}$) $s=\log(\alpha)\times(1/(\sqrt{2}\theta_{1})+0.088)$ and $c=-\log(1-\sqrt{1-\alpha})/(2\theta_{1})-0.583$ with $\alpha=0.05$. Table 6.11 reports the average size and coverage rates (for target conditional coverage 0.950.950.95) across 500500500 runs.
Table 6.11: Comparison of the confidence sets for $T$ with [36] (Setting I)
$F_{0}$ $F_{1}$ $T$ Conditional Size Conditional Coverage Marginal Coverage
Ours Wu (2007) Ours Wu (2007) Ours Wu (2007)
${N}(-0.25,1)$ ${N}(0.25,1)$ 100 46.39 47.83 0.96 0.86 0.88 0.79
${N}(-0.3,1)$ ${N}(0.3,1)$ 100 33.74 34.14 0.96 0.83 0.82 0.76
${N}(-0.25,1)$ ${N}(0.25,1)$ 500 46.58 47.21 0.96 0.86 0.57 0.53
${N}(-0.3,1)$ ${N}(0.3,1)$ 500 35.19 34.81 0.97 0.82 0.56 0.49
The results show that [36] (which is only asymptotically valid and relies on approximations) exhibits noticeable undercoverage, while the sizes for both methods are comparable. However, our method (4.1) achieves the target conditional coverage.